# Test-time Adaptive Vision-and-Language Navigation

### v2

@@ 3.2 Fast Update via Gradient Analysis
At timestep $t$ in the navigation process, the agent is required to select an action (navigable node) by using the predicted score ${\bm{s}}_{t}$. With this score, we can calculate the TTA loss (Eq. (2)) and then derive the gradient of the model parameters $\bm{\Theta}$ as: $\bm{g}_{t}=\nabla\mathcal{L}({\bm{s}}_{t};\bm{\Theta})$, $\bm{g}_{t}\in\mathbb{R}^{D}$. Traditional TTA methods conduct adaptation independently at each time step, which can exacerbate the issue of cumulative errors [50, 60], particularly in the VLN process that requires frequent action execution. Therefore, we propose to conduct a gradient decomposition-accumulation analysis, wherein we periodically analyze the gradients generated during the recent multi-step navigation process and identify a concordant direction for an iteration of model update.
Gradient Decomposition-Accumulation. During navigation, as shown in Figure 2, we perform model update every $M$ action steps. For the $j$-th update, gradients from previous $M$ steps are collected as $\bm{G}_{j}=\{\widetilde{\bm{g}}_{j,m}\}_{m=1}^{M}$, where $\bm{G}_{j}\in\mathbb{R}^{M\times D}$, $\widetilde{\bm{g}}_{j,m}$ indicates the $t$-th gradient $\bm{g}_{t}$ when $t=M(j-1)+m$. Note that these gradients determines the learning direction of our VLN model, and a simple strategy to compute this direction is to take their average $\bar{\bm{g}}_{j}=1/M\sum_{m}\widetilde{\bm{g}}_{j,m}$; however, this inevitably introduce step-specific noise. To avoid the issue, we aim to find a concordant direction among these gradients. We first establish a local coordinate system with $D$ orthogonal and unit axes (bases) $\bm{U}_{j}=\{{\bm{u}_{j,d}}^{\mathsf{T}}\}_{d=1}^{D}\in\mathbb{R}^{D\times D}$ for gradient decomposition, where each gradient can be approximatively linearly represented by these bases. Intuitively, the axes along which the gradients exhibit the higher variance after projection represent the directions of gradients with the lower consistency. These directions have the potential to introduce interference in determining a model update direction. Therefore, it is advisable to reduce the projection of gradients in these directions. To solve the bases $\bm{U}_{j}$, we can utilize singular value decomposition (SVD) as follows:
@@
where $\hat{\bm{G}}_{j}$ is the centered gradient matrix by removing the mean from $\bm{G}_{j}$. The $m$-th row vector in $\hat{\bm{G}}_{j}$ reflects the deviation between $\widetilde{\bm{g}}_{j,m}$ and the average gradient $\bar{\bm{g}}_{j}$. $\lambda_{j,d}$, $\bm{u}_{j,d}$ denote the $d$-th largest eigenvalue and the corresponding eigenvector. Motivated by the principle component analysis [59], it is obvious that a larger $\lambda_{j,d}$ corresponds to a higher variance of the gradient projection length $\bm{G}_{j}\bm{u}_{j,d}$ and vice versa. Hence, we can derive a concordant gradient by adaptively aggregating the gradients’ components on all the axes by considering different eigenvalue (importance):
@@
With $\nabla^{(fast)}$, we can perform fast model update by setting a learning rate $\gamma^{(fast)}$. Although traditional methods employ a fixed learning rate during optimization, such a setting might hinder model convergence, i.e., small learning rates slow down convergence while aggressive learning rates prohibit convergence [4]. Since fast updates are frequently invoked during navigation, relying on a fixed learning rate is sub-optimal. Therefore, we propose to dynamically adjust learning rate throughout the fast update phase.
Dynamic Learning Rate Scaling. Different from varying the learning rate through optimizer or scheduler, we argue for a scaling method that leverages gradient agreement information in historical steps to dynamically adjust the speed of model update. Current gradient alignment strategies typically impose direct constraints on the gradients [58, 56], which are not suitable for our framework as they undermine the gradient decomposition-accumulation process. Given that the second-order information (variance) has been demonstrated to be more effective than the first-order information (mean) in gradient agreement learning [56], we directly utilize the trace of the gradient covariance matrix, $\text{Tr}\left(1/(M-1){\hat{\bm{G}}}_{j}^{\mathsf{T}}\hat{\bm{G}}_{j}\right)$, for scaling. Note that the trace is equal to the sum of eigenvalues $\sigma_{j}=\sum_{d}\lambda_{j,d}$. Here, when $\sigma_{j}$ deviates significantly from the historical variance, we assign a smaller learning rate, and vice versa:
@@
Model Update. With the above gradient and learning rate, we can perform the $j$-th fast model update:
@@ 3.3 Slow Update via Parameter Analysis
In the fast update phase, although we obtain concordant optimization directions, the frequent parameter updates may still dramatically change the VLN model. To maintain the stability of the VLN model during long-term usage, we revert the model to its historical states recorded in the fast update phase, and conduct a decomposition-accumulation analysis on the parameter variation trajectory for direct parameter modulation. The slow update phase shares the core formulation with the fast phase, but shifts the focus from gradients to the model parameters themselves.
@@
Table 1: Comparison on REVERIE. Results better than the base model are highlighted in bold.
Table 2: Results on R2R dataset.
Methods REVERIE Val Unseen REVERIE Test Unseen
TL $\downarrow$ OSR SR SPL RGS RGSPL TL $\downarrow$ OSR SR SPL RGS RGSPL
Human - - - - - - 21.18 86.83 81.51 53.66 77.84 51.44
Seq2Seq [3] [CVPR18] 11.07 8.07 4.20 2.84 2.16 1.63 10.89 6.88 3.99 3.09 2.00 1.58
RCM [74] [CVPR19] 11.98 14.23 9.29 6.97 4.89 3.89 10.60 11.68 7.84 6.67 3.67 3.14
SMNA [45] [ICLR19] 9.07 11.28 8.15 6.44 4.54 3.61 9.23 8.39 5.80 4.53 3.10 2.39
FAST [52] [CVPR20] 45.28 28.20 14.40 7.19 7.84 4.67 39.05 30.63 19.88 11.61 11.28 6.08
Airbert [21] [ICCV21] 18.71 34.51 27.89 21.88 18.23 14.18 17.91 34.20 30.28 23.61 16.83 13.28
HAMT [9] [NeurIPS21] 14.08 36.84 32.95 30.20 18.92 17.28 13.62 33.41 30.40 26.67 14.88 13.08
HOP [53] [CVPR22] 16.46 36.24 31.78 26.11 18.85 15.73 16.38 33.06 30.17 24.34 17.69 14.34
LANA [75] [CVPR23] 23.18 52.97 48.31 33.86 32.86 22.77 18.83 57.20 51.72 36.45 32.95 22.85
BEVBert [1] [ICCV23] - 56.40 51.78 36.37 34.71 24.44 - 57.26 52.81 36.41 32.06 22.09
BSG [44] [ICCV23] 24.71 58.05 52.12 35.59 35.36 24.24 22.90 62.83 56.45 38.70 33.15 22.34
GridMM [78] [ICCV23] 23.20 57.48 51.37 36.47 34.57 24.56 19.97 59.55 55.13 36.60 34.87 23.45
DUET [11] [CVPR22] 22.11 51.07 46.98 33.73 32.15 23.03 21.30 56.91 52.51 36.06 31.88 22.06
DUET-FSTTA 22.14 56.26 54.15 36.41 34.27 23.56 21.52 58.44 53.40 36.43 32.99 22.40
HM3D [10] [ECCV22] 22.13 62.11 55.89 40.85 36.58 26.76 20.87 59.81 53.13 38.24 32.69 22.68
HM3D-FSTTA 22.37 63.74 57.02 41.41 36.97 26.55 21.90 63.68 56.44 39.58 34.05 23.04
Methods R2R Val Unseen
TL $\downarrow$ NE $\downarrow$ SR SPL
Seq2Seq [3] 8.39 7.81 22 -
RCM [74] 11.46 6.09 43 -
SMNA [45] - 5.52 45 32
EnvDrop [63] 10.70 5.22 52 48
AirBert [21] 11.78 4.10 62 56
HAMT [9] 11.46 3.65 66 61
GBE [86] - 5.20 54 43
SEvol [7] 12.26 3.99 62 57
HOP [53] 12.27 3.80 64 57
BEVBert [1] 14.55 2.81 75 64
LANA [75] 12.00 - 68 62
BSG [44] 14.90 2.89 74 62
GridMM [78] 13.27 2.83 75 64
DUET [11] 13.94 3.31 72 60
DUET-FSTTA 14.64 3.03 75 62
HM3D [10] 14.29 2.83 74 62
HM3D-FSTTA 14.86 2.71 75 63
Table 2: Results on R2R dataset.
Similar to the fast update phase, the centered parameter matrix $\hat{\bm{M}}_{l}$ can be constructed, where the $n$-th row vector in it reflects the deviation between $\widetilde{\bm{\Theta}}_{l,n}$ and the averaged historical parameter $\bar{\bm{\Theta}}_{l}=1/(N+1)\sum_{n}\widetilde{\bm{\Theta}}_{l,n}$. With $\hat{\bm{M}}_{l}$, we can obtain the following eigenvalues and eigenvectors: $\epsilon_{l,d},\bm{z}_{l,d}=\textbf{SVD}_{d}(1/N\cdot{\hat{\bm{M}}}_{l}^{\mathsf{T}}\hat{\bm{M}}_{l})$, where a larger $\epsilon_{l,d}$ corresponds to a higher variance of the parameter projection length ${\bm{M}}_{l}\bm{z}_{l,d}$ and vice versa. $\bm{Z}_{l}=\{{\bm{z}_{l,d}}^{\mathsf{T}}\}_{d=1}^{D}$ depicts the local coordinate system where each axis depicts the direction of parameter variation. Intuitively, the principal axes (with larger eigenvalues) delineate the primary directions of historical parameter variation, while minor axes (with smaller eigenvalues) often encompass noise [76]. To find a more reliable optimization path to traverse the trajectory of primary parameter changes, we pay more attention on the axis with the larger variance. Since there is no silver bullet to learning an optimization direction with only parameters, a reference direction can significantly aid in guiding the model towards a local optimal. Here, we leverage the parameter variations to calculate the reference direction:
where the L2-normalization is performed on eigenvalues to convey different relative importance of axes. Besides, the norm of the reference direction is utilized to automatically tuning the magnitude of the analyzed gradient. In contrast to $\Phi_{d}(\cdot)$ in gradient analysis, $\Psi_{d}(\cdot)$ highlights those axes with high variation due to the different characteristics of gradients and parameters in model optimization.
where $\gamma^{(slow)}$ is learning rate. Since the slow update phase is designed for stable model learning and is not frequently invoked, we employ a fixed learning rate here instead of conducting dynamic learning rate scaling, as done in the fast phase. The updated parameter $\bm{\Theta}^{(l)}$ will be utilized for the subsequent test samples in conjunction with new fast update phases applied to them.
4 Experimental Results
We evaluate FSTTA on four benchmarks: REVERIE [52], R2R [3], SOON [86], and R2R-CE [32] datasets. Experiments and ablation studies show our effectiveness.
4.1 Experimental Setup
Datasets. Four datasets are adopted for conducting our experiments. Among them, REVERIE [52] contains 10,567 panoramic images and 21,702 high-level instructions, focusing on grounding remote target object within 90 buildings. R2R [3] provides step-by-step instructions for navigation in photo-realistic environments, which includes 10,800 panoramic views and 7,189 trajectories. SOON [86] also requires the agent to find the target object with a more detailed description of the goal. It has 3,848 sets of instruction and more than 30K long distance trajectories. R2R-CE [32] is a variant of R2R in continuous environments, where an agent is able to move freely and engage with obstacles. The dataset consists of 16,000 instruction-trajectory pairs, with non-transferrable paths excluded.
Evaluation Metrics. We follow previous approaches [52, 11, 10, 37, 78] and employ the most commonly used metrics for evaluating VLN agents as follows: TL (Trajectory Length), NE (Navigation Error), SR (Success Rate), SPL (Success weighted by Path Length), OSR (Oracle Success Rate), RGS (Remote Grounding Success rate), and RGSPL (RGS weighted by Path Length).
Implementation Details. To better conform to practical scenarios, for all datasets, we set the batch size to 1 during evaluation. Each sample (or each action step) is forward propagated only once during the testing process. We adopt DUET [11] and HM3D [10] as the base models. Since HM3D does not provide training code for R2R-CE dataset, we adopt another state-of-the-art method, BEVBert [1], for TTA. Note that for the base models, in Section 4.2, we report the results obtained from running their official codes. For VLN models equipped with TTA strategies, we run the corresponding experiments 5 times while shuffling the order of the samples and report the average results. In our FSTTA, we only utilize the last four LN layers of base models for model updating, all the feature dimensions of these layers are 768. We set the intervals for fast and slow updates to $M=3$ and $N=4$, the learning rates of the two phases are $\hat{\gamma}^{(fast)}=6\times 10^{-4}$ and $\gamma^{(slow)}=1\times 10^{-3}$. For the dynamic learning rate scaling, we empirically set the threshold $\tau=0.7$ in Eq. (6) and the update momentum $\rho=0.95$ with the truncation interval $[0.9,1.1]$. And the hyper-parameter $q$ in Eq. (8) is set to 0.1. All experiments are conducted on a RTX 3090 GPU.
4.2 Comparison with State-of-the-art VLN Models
REVERIE. Table 2 presents a comparison of our FSTTA against state-of-the-art methods on the REVERIE dataset. Compared with the base models which do not perform test-time adaptation, the proposed method demonstrates favorable performance improvement across most evaluation metrics across the two dataset splits. Specifically, on the validation unseen split, our model exhibits notable advantages over DUET, with improvements of 5.3% on OSR, 7.1% on SR, and 2.7% on SPL. Furthermore, for the recent state-of-the-art method HM3D, our model displays enhanced generalization capabilities on the test unseen split, achieving remarkable improvements over HM3D, including 3.9%, 3.3%, and 1.3% increases on the three metrics. Compared with other state-of-the-arts, our proposed method can achieve superior or comparable performance. These results unequivocally affirm the effectiveness of our fast-slow test time adaptation model, showing the promising potential of TTA in the VLN field. It is noteworthy that none of the prior methods employed a TTA strategy on this task.
R2R. Table 2 shows the comparison results on R2R dataset. Our approach outperforms the base models in most metrics (e.g., $72\%\rightarrow 75\%$ for DUET on SR, $62\%\rightarrow 63\%$ for HM3D on SPL). Notably, from the results of the above two datasets, our method, while enhancing the success rate of VLN, causes a slight increase in the path length (TL). We speculate that a possible reason is that performing TTA online may increase the likelihood of the agent deviating from its original action execution pattern, leading to more exploration or backtracking. This situation is further confirmed in the analysis of various TTA strategies in Table 5.
Table 3: Experimental results on SOON dataset.
Methods Val Unseen Test Unseen
OSR SR SPL RGSPL OSR SR SPL RGSPL
GBE [86] 28.54 19.52 13.34 1.16 21.45 12.90 9.23 0.45
GridMM [78] 53.39 37.46 24.81 3.91 48.02 36.27 21.25 4.15
DUET [11] 50.91 36.28 22.58 3.75 43.00 33.44 21.42 4.17
DUET-FSTTA 52.57 36.53 23.82 3.75 43.44 35.34 23.23 4.52
HM3D [10] 53.22 41.00 30.69 4.06 47.26 40.26 28.09 5.15
HM3D-FSTTA 54.19 42.44 31.03 4.93 48.52 42.02 28.95 5.20
R2R-CE. FSTTA also generalizes well on the continuous environment, i.e., R2R-CE dataset, as shown in Table 4. The results indicate that our approach demonstrates superior or comparable performance against other methods across several metrics.
Table 4: Experimental results on R2R-CE dataset.
Methods Val Unseen Test Unseen
NE $\downarrow$ OSR SR SPL NE $\downarrow$ OSR SR SPL
Seq2Seq [32] 7.37 40 32 30 7.91 36 28 25
CWTP [8] 7.90 38 26 23 - - - -
[…]
Sim2Sim [31] 6.07 52 43 36 6.17 52 44 37
CWP-BERT [26] 5.74 53 44 39 5.89 51 42 36
DREAMW [69] 5.53 49 59 44 5.48 49 57 44
GridMM [78] 5.11 61 49 41 5.64 56 46 39
ETPNav [2] 4.71 65 57 49 5.12 63 55 48
DUET [11] 5.13 55 46 40 5.82 50 42 36
DUET-FSTTA 5.27 58 48 42 5.84 55 46 38
BEVBert [1] 4.57 67 59 50 4.70 67 59 50
BEVBert-FSTTA 4.39 65 60 51 5.45 69 60 50
Table 5: Experimental results for different TTA strategies.
Methods REVERIE Val Unseen Time(ms)
TL $\downarrow$ OSR SR SPL RGS RGSPL
DUET [11] 22.11 51.07 46.98 33.73 32.15 23.03 104.84
 + EATA [50] 23.41 52.09 47.40 33.46 32.09 22.65 133.12
[…]
 + NOTE [19] 23.15 52.85 48.28 33.98 32.77 22.98 137.89
 + SAR [51] 23.47 53.26 48.00 33.92 33.49 23.09 145.53
 + Tent [68] 24.05 49.43 46.87 31.90 30.04 20.15 126.91
 + Tent-INT-2 24.24 51.22 48.46 33.67 32.43 21.30 124.02
 + Tent-INT-3 22.52 52.28 48.60 34.65 32.66 23.12 119.34
 + Tent-INT-4 22.59 51.40 48.91 35.06 32.59 22.99 117.26
 + Tent-Stable 22.05 51.43 47.55 33.99 32.34 23.32 129.22
 + FSTTA 22.14 56.26 54.15 36.41 34.27 23.56 135.61
4.3 Results for Different TTA Strategies
Currently, various TTA methods have been adeptly integrated for the dynamic model updates within diverse computer vision tasks, marking significant progress. Although the exploration of TTA’s application within the VLN field remains relatively untapped, the integration of contemporary advanced TTA methodologies into VLN is feasible. Since efficiency is an important evaluation metric for TTA, we provide the average time taken by each method to execute a single instruction for comparison. Obviously, equipping with TTA inevitably incurs additional time costs. For the compared methods, SAR and TENT are the popular entropy minimization models, whereas NOTE, CoTTA, and EATA are state-of-the-art continual TTA methods. The results in Table 5 demonstrate the capability of our proposed FSTTA to blend model performance with testing efficiency. Specifically, on the validation unseen dataset of REVERIE, our method exhibits a discernible enhancement of 6.2% and 2.5% on the SR and SPL metrics compared to the state-of-the-art SAR method, concurrently manifesting a reduction of 7% in testing time. From the results, directly applying existing TTA methods to the VLN task does not lead to significant performance improvements. Furthermore, we investigate different frequencies of updates based on TENT as well as the stable update approach. ‘INT’ represents the update interval, which means averaging the gradient information over a certain interval and then performing an iteration of model update; these results are consistent with those in Figure 1(b). It can be seen that our method still outperforms these strategies with marginally increased time costs.
@@ 4.3 Results for Different TTA Strategies
- - 13.86 73.86 71.15 63.94 57.41 51.14
- 15.13 75.59 75.48 65.84 58.62 52.23
- 13.40 73.16 71.78 64.18 57.05 51.18
15.11 75.58 74.12 65.53 59.20 52.18
4.4 Further Remarks
We perform ablation studies and other in-depth analysis of FSTTA on the validation unseen set of REVERIE [52].
Ablation Studies of the Proposed FSTTA. In this work, we propose a FSTTA method for vision-and-language navigation, which consists of both fast and slow model update phases. To validate their effectiveness, we progressively integrate the two phases into the baseline DUET model. In addition, we design a baseline variant, which equips DUET with the vanilla TTA objective (TENT [68]) and simply utilize the averaged gradient in an interval (with the same $M$) for fast model updates. Empirical findings from Table 6 illuminate that the integration of fast and slow phases progressively bolsters the base model by 2.8% and 4.3% on the SR metric. Moreover, the dynamic learning rate scaling module (DLR) also contributes to enhancing the model’s performance. Furthermore, our method surpasses the vallina TTA method by a significant margin, showing the consideration of the fast-slow update mechanism is effective.
Will our method experience catastrophic forgetting? For a VLN agent endowed with the TTA capability, it faces the issue of catastrophic forgetting of historical environments and instructions upon continually executing new instructions in new environments. To assess whether our method harbors this issue, we re-evaluate our methods on REVERIE validation seen data. Compared with the base model, as shown in Table 7, we find that: (1) Directly applying FSTTA with the base model on seen data can noticeably enhance performance. (2) After performing FSTTA on the unseen set, the obtained model, when tested directly on the seen dataset without TTA, achieves performance comparable to the base model, confirming that our method does not suffer from catastrophic forgetting. (3) Applying the updated model from unseen set to the seen set with TTA yields the best results. This indicates that our method is effective in accumulating experience from historical test data.
Generalization Testing in More Practical Environments. In the real-world applications, agents might encounter both previously seen and unseen scenarios. In our preceding experiments, we exclusively test on the validation seen and unseen sets separately. To verify the generalizability, we combine the seen and unseen sets into a unified set. Table 8 shows that FSTTA outperforms other TTA methods in effectively managing a variety of testing scenarios.
Table 8: Results of validation unseen & seen on REVERIE dataset.

### v3

@@ 3.1 Fast Update via Gradient Analysis
At timestep $t$ in one navigation process, the agent is required to select an action (navigable node) by using the predicted score ${\bm{s}}_{t}$. With ${\bm{s}}_{t}$, we can calculate the TTA loss (Eq. (2)) and then derive the gradient of the model parameters $\bm{\Theta}$ as: $\bm{g}_{t}=\nabla\mathcal{L}({\bm{s}}_{t};\bm{\Theta})$, $\bm{g}_{t}\in\mathbb{R}^{D}$. Traditional TTA methods conduct adaptation independently at each time step, which can exacerbate the issue of cumulative errors (Niu et al., 2022; Song et al., 2023), particularly in the VLN process that requires frequent action execution. Therefore, we propose to conduct a gradient decomposition-accumulation analysis, wherein we periodically analyze the gradients generated during the recent multi-step navigation process and identify a concordant direction for an iteration of model update.
Gradient Decomposition-Accumulation. During navigation, as shown in Figure 2, we perform model update every $M$ action steps. For the $j$-th update, gradients from previous $M$ steps are collected as $\bm{G}_{j}=\{\widetilde{\bm{g}}_{j,m}\}_{m=1}^{M}$, where $\bm{G}_{j}\in\mathbb{R}^{M\times D}$, $\widetilde{\bm{g}}_{j,m}$ indicates the $t$-th gradient $\bm{g}_{t}$ when $t=M(j-1)+m$. Note that these gradients determine the learning direction of our VLN model, and a simple strategy to compute this direction is to take their average $\bar{\bm{g}}_{j}=1/M\sum_{m}\widetilde{\bm{g}}_{j,m}$; however, this inevitably introduce step-specific noise. To avoid the issue, we aim to find a concordant direction among these gradients. We first establish a local coordinate system with $D$ orthogonal and unit axes (bases) $\bm{U}_{j}=\{{\bm{u}_{j,d}}^{\mathsf{T}}\}_{d=1}^{D}\in\mathbb{R}^{D\times D}$ for gradient decomposition, where each gradient can be approximatively linearly represented by these bases. Intuitively, the axes along which the gradients exhibit the higher variance after projection represent the directions of gradients with the lower consistency. These directions have the potential to introduce interference in determining a model update direction. Therefore, it is advisable to reduce the projection of gradients in these directions. To solve the bases $\bm{U}_{j}$, we can utilize singular value decomposition (SVD) as follows:
@@
where $\hat{\bm{G}}_{j}$ is the centered gradient matrix by removing the mean from $\bm{G}_{j}$. The $m$-th row in $\hat{\bm{G}}_{j}$ reflects the deviation between $\widetilde{\bm{g}}_{j,m}$ and the average gradient $\bar{\bm{g}}_{j}$. $\lambda_{j,d}$, $\bm{u}_{j,d}$ denote the $d$-th largest eigenvalue and the corresponding eigenvector. Motivated by the principle component analysis (Shlens, 2014), it is obvious that a larger $\lambda_{j,d}$ corresponds to a higher variance of the gradient projection length $\bm{G}_{j}\bm{u}_{j,d}$ and vice versa. Hence, we can derive a concordant gradient by adaptively aggregating the gradients’ components on all the axes by considering different eigenvalue (importance):
@@
With $\nabla^{(fast)}$, we can perform fast model update by setting a learning rate $\gamma^{(fast)}$. Although traditional methods employ a fixed learning rate, such a setting might hinder model convergence, i.e., small learning rates slow down convergence while aggressive learning rates prohibit convergence (Barzilai & Borwein, 1988). Since fast updates are frequently invoked during navigation, relying on a fixed learning rate is sub-optimal. Therefore, we propose to dynamically adjust learning rate throughout the fast update phase.
Dynamic Learning Rate Scaling. Different from varying the learning rate through optimizer or scheduler, we argue for a scaling method that leverages gradient agreement information in historical steps to dynamically adjust the speed of model update. Current gradient alignment strategies typically impose direct constraints on the gradients (Shi et al., 2021; Rame et al., 2022), which are not suitable for our framework as they undermine the gradient decomposition-accumulation process. Given that the second-order information (variance) has been demonstrated to be more effective than the first-order information (mean) in gradient agreement learning (Rame et al., 2022), we directly utilize the trace of the gradient covariance matrix, $\text{Tr}\left(1/(M-1){\hat{\bm{G}}}_{j}^{\mathsf{T}}\hat{\bm{G}}_{j}\right)$, for scaling. Note that the trace is equal to the sum of eigenvalues $\sigma_{j}=\sum_{d}\lambda_{j,d}$. Here, when $\sigma_{j}$ deviates significantly from the historical variance, we assign a smaller learning rate, and vice versa:
@@
Model Update. With the above gradient and learning rate, we can perform a single iteration as the $j$-th fast update:
@@ 3.2 Slow Update via Parameter Analysis
In the fast update phase, although we obtain concordant optimization directions, the frequent parameter updates may still dramatically change the VLN model. To maintain the stability of the VLN model during online long-term usage, we periodically revert the model to its historical states, and conduct a decomposition-accumulation analysis on the parameter variation trajectory for direct parameter modulation. The slow update phase shares the core formulation with the fast phase, but shifts the focus from gradients to the model parameters themselves.
@@
Table 1: Experimental results for different TTA strategies on REVERIE dataset.
Methods REVERIE Val Seen REVERIE Val Unseen REVERIE Test Unseen Time(ms)*
OSR SR SPL RGSPL OSR SR SPL RGSPL OSR SR SPL RGSPL
DUET (Chen et al., 2022d) 73.86 71.75 63.94 51.14 51.07 46.98 33.73 23.03 56.91 52.51 36.06 22.06 104.84
   + EATA (Niu et al., 2022) 74.05 72.17 63.41 49.94 52.09 47.40 33.46 22.65 57.21 52.91 35.19 21.65 133.12
   + CoTTA (Wang et al., 2022a) 73.79 71.05 62.91 49.36 52.46 47.56 31.43 21.83 56.14 52.52 34.66 21.06 3.89$\times$ 103
   + NOTE (Gong et al., 2022) 74.76 72.43 64.28 51.97 52.85 48.28 33.98 22.98 57.66 53.41 36.18 22.09 137.89
   + SAR (Niu et al., 2023) 74.84 71.75 64.43 51.70 53.26 48.00 33.92 23.09 57.11 53.04 36.07 22.27 145.53
   + ViDA (Liu et al., 2024) 73.99 72.49 63.49 50.89 52.53 48.14 32.45 21.92 56.78 52.74 35.10 21.77 3.97$\times$ 103
   + Tent (Wang et al., 2021) 70.07 70.73 61.67 49.31 49.43 46.87 31.90 20.15 54.58 50.56 33.37 20.32 126.91
   + Tent-INT-2 71.82 70.36 61.53 49.82 51.22 48.46 33.67 21.30 54.25 50.36 33.89 21.09 124.02
   + Tent-INT-3 74.54 72.58 64.44 51.05 52.28 48.60 34.65 23.12 56.66 52.92 35.84 21.89 119.34
   + Tent-INT-4 73.84 72.51 64.03 50.97 51.40 48.91 35.06 22.99 56.74 53.14 36.29 22.14 117.26
   + Tent-Stable 73.72 71.89 64.06 50.41 51.43 47.55 33.99 23.32 57.12 52.61 36.17 22.16 129.22
   + FSTTA 75.59 75.48 65.84 52.23 56.26 54.15 36.41 23.56 58.44 53.40 36.43 22.40 135.61
    Note that the last column displays the average execution time of the agent for a single instruction, calculated on the Validation Unseen set of REVERIE.
@@
 $\times$           $\times$ 13.86 73.86 71.15 63.94 57.41 51.14
 $\times$  $\longrightarrow$  ✓ 15.13 75.59 75.48 65.84 58.62 52.23
 ✓  $\longrightarrow$  $\times$ 13.40 73.16 71.78 64.18 57.05 51.18
 ✓  $\longrightarrow$  ✓ 15.11 75.58 74.12 65.53 59.20 52.18
Similar to the fast update phase, the centered parameter matrix $\hat{\bm{M}}_{l}$ can be constructed, where the $n$-th row vector in it reflects the deviation between $\widetilde{\bm{\Theta}}_{l,n}$ and the averaged historical parameter $\bar{\bm{\Theta}}_{l}=1/(N+1)\sum_{n}\widetilde{\bm{\Theta}}_{l,n}$. With $\hat{\bm{M}}_{l}$, we can obtain the following eigenvalues and eigenvectors: $\epsilon_{l,d},\bm{z}_{l,d}=\textbf{SVD}_{d}(1/N\cdot{\hat{\bm{M}}}_{l}^{\mathsf{T}}\hat{\bm{M}}_{l})$, where a larger $\epsilon_{l,d}$ corresponds to a higher variance of the parameter projection length ${\bm{M}}_{l}\bm{z}_{l,d}$ and vice versa. $\bm{Z}_{l}=\{{\bm{z}_{l,d}}^{\mathsf{T}}\}_{d=1}^{D}$ depicts the local coordinate system where each axis depicts the direction of parameter variation. Intuitively, the principal axes (with larger eigenvalues) delineate the primary directions of historical parameter variation, while minor axes (with smaller eigenvalues) often encompass noise (Wang et al., 2023e). To find a reliable optimization path to traverse the trajectory of primary parameter changes, we pay attention on the axis with the large variance. Since there is no silver bullet to learning an optimization direction with only parameters, a reference direction can significantly aid in guiding the model towards a local optimal. Here, we leverage the parameter variations to calculate the reference direction:
where the L2-normalization is performed on eigenvalues to convey different relative importance of axes. Besides, the norm of the reference direction is utilized to automatically tuning the magnitude of the analyzed gradient. In contrast to $\Phi_{d}(\cdot)$ in the fast update phase, $\Psi_{d}(\cdot)$ highlights those axes with high variation due to the different characteristics of gradients and parameters in model optimization.
where $\gamma^{(slow)}$ is learning rate. Since the slow update phase is designed for stable model learning and is not frequently invoked, we employ a fixed learning rate here instead of conducting the dynamic learning rate scaling as done in the fast phase. The updated parameter $\bm{\Theta}^{(l)}$ will be utilized for the subsequently coming test samples in conjunction with new fast update phases applied to them.
4 Experimental Results
Table 4: Results on Validation Unseen & Seen sets of REVERIE.